OpenStudy (anonymous):

Osmosis can be proved by Molecular weight. how would we be able to prove Diffusion?

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

Diffusion is the spread of one substance throughout a space or throughout a second substance. If a molecule is diffusing, then it would be proved by its molecular speed.

OpenStudy (anonymous):

The faster molecular speed, the higher the diffusion is.

OpenStudy (anonymous):

Got it thanks a lot Shalante :)

OpenStudy (anonymous):

No problem!

OpenStudy (jhannybean):

Along with what @Shalante mentioned, \[\sf KE ~~\alpha~ \frac{1}{molecular ~weight}\]\[\sf KE ~\alpha~ T^\circ\]

OpenStudy (greatlife44):

There's another law too don't know if it directly applies here but it's called Graham's law of effusion/diffusion Something like this R1/R2 = (m2/m1)^(1/2) The rate of diffusion of a gas is inversely proportional to the mass of that gas. The heavier the gas the slower it effuses/diffuses.